OpenStudy (anonymous):

Suppose a triangle has sides a, b, and c, and let be the angle opposite the side of length b. If cos > 0, what must be true? A. a^2 + c^2 < b^2 B. a^2 + b^2 = c^2 C. a^2 + c^2 > b^2 D. a^2 + b^2 > c^2 **i don't get this one :( @terenzreignz :)

terenzreignz (terenzreignz):

Well, remember that \[\large \cos(\theta) = \frac{a^2+c^2-b^2}{2ac}\]

OpenStudy (anonymous):

mhmm :)

terenzreignz (terenzreignz):

Well then, since a, b, and c are all positive, then 2ac is positive, so it all falls on a^2 + c^2 - b^2

OpenStudy (anonymous):

yes :) so would that mean a^2 + c^2 > b^2 ??

OpenStudy (anonymous):

is that right? :)